0.3x-0.2y=4
0.5x+0.5y=-45/17
simplified, to include ordered pair

Answer by ewatrrr(24785)

 
Hi
0.3x-0.2y=4
0.5x+0.5y=-45/17
0.3x-0.2y=4
x + y =-90/17
0.3x-0.2y =4
0.2x +0.2y =-90/17*2/10 = -18/17
.5x = 50/17
x = 100/17 and y = -190/17 .5y = -95/17
CHECKING our Answer***
3/10*100/17 -2/10*-190/17 = 30/17 + 38/17 = 68/17 = 4